In the right triangle the ratio of the measures of three sides are 2 4 8, and its perimeter is 280 inches. find the length to the longest side.
Found 2 solutions by mananth, richard1234:
Answer by mananth(16949)

the sides are in the ration 2:4:8
let the common multiple be x
the sides are 2x , 4x,8x
perimeter = 280
2x+4x+8x=280
14x=280
/14
x= 20
The length of the largest side = 8x = 8*20 =160 in

Answer by richard1234(7193)

A triangle with sides 2x, 4x, 8x can't be a triangle (this violates the triangle inequality).